OpenStudy (anonymous):

How far is point P from point Q? assessment graphic A. 7 units B. 9 units C. 10 units D. 8 units

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@demonchild99

OpenStudy (anonymous):

@jeffyblood

OpenStudy (anonymous):

@HAIIll

OpenStudy (anonymous):

@skate-4-life270

OpenStudy (anonymous):

Just count the lines!

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

yes 7

OpenStudy (anonymous):

i know but i am confuse the middle the one that says not number ?!?!?!?!?!

OpenStudy (jeffyblood):

@trinajoy do you need my help at all any more??

OpenStudy (anonymous):

What is the perimeter of this rectangular field?

OpenStudy (jeffyblood):

Multiply the numbers

OpenStudy (anonymous):

do (78.6)^2+(93.2)^2

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Not muliply... thats area

OpenStudy (anonymous):

7

OpenStudy (anonymous):

(78.6*2)+(93.2*2)

OpenStudy (anonymous):

Do that

OpenStudy (jeffyblood):

Sorry I was thinking that it said area but I am just going crazy lol

OpenStudy (anonymous):

lol @jeffyblood

OpenStudy (anonymous):

343.6

OpenStudy (anonymous):

is it 343.6? @JohnKnows

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Anne walked 3.2 kilometers. How many meters did Anne walk? A. 0.032 meters B. 3200 meters C. 0.0032 meters D. 320 meters

OpenStudy (anonymous):

Please please i need help @JohnKnows @jeffyblood

OpenStudy (anonymous):

3200

OpenStudy (anonymous):

B

OpenStudy (jeffyblood):

Well there is 0.001 in one meter

OpenStudy (anonymous):

so it is?

OpenStudy (jeffyblood):

B? yes

OpenStudy (anonymous):

3,200

OpenStudy (anonymous):

Every Km is 1,000 meters

OpenStudy (jeffyblood):

@JohnKnows yes that means that it is B

OpenStudy (anonymous):

yep :) @jeffyblood

OpenStudy (anonymous):

he perimeter of this rectangle can be calculated using the equation P = 2 × 18.5 + 2 × 7.2. Which other equation could be used to calculate the perimeter? A. P = (2 • 18.5) + 7.2 B. P = 2 (18.5 + 7.2) C. P = 2 × 18.5 + 7.2 D. P = 2 × 18.5 × 7.2

OpenStudy (anonymous):

C